Solve by factoring and using the principle of zero products. +2 - X = 20 O 1, 20 4,5 4,5 -4,5
Solve by factoring and using the principle of X -- X =
Answer #1

922_0 = 20 72_9 -~-20 = 36-20 42_4-90=0 3 - 511 4 3 – 2 = 0 a CX-5) + 4(x-5)=0 (U+4)(0-5)=0 from the principle of zero produc